Circular No. 7785

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

PECULIAR VARIABLE IN MONOCEROS N. J. Brown, Quinns Rocks, W. Australia, reports his discovery of a

possible nova on a T-Max 400 film taken on Jan. 6.6 UT (mag ∼ 10), confirmed visually at mag 10.0 on Jan. 7.6. Brown noted the object to be located at α = 7^{h}04^{m}.0, δ = −3^{o}50′ (equinox 2000.0), adding that nothing appeared at this position (limiting mag 12) on a photo taken on 2001 Dec.

22. E. O. Waagen, AAVSO, reports confirming observations by C. Scovil (Stamford, CT; CCD mag V = 10.1 on Jan. 8.094), by P. Nelson (Ellinbank, Victoria; unfiltered CCD mag 9.3 on Jan. 9.513), and by A. Oksanan and J. Solonen (Nyrölä, Finland; unfiltered CCD mag 9.5 on Jan. 9.815). Oksanan provides the following precise position and photometry for the star on Jan. 9.96: α = 7^{h}04^{m}04^{s}.85, δ = −3^{o}50′51′′.1, V = 9.92, B − V = +1.72 (around 30 USNO-A2.0 reference stars; Bessel filters; Tycho-catalogue reference stars for photometry). A. Price, AAVSO, writes that a search of 71 RH photographic plates at Harvard College Observatory from 1930 Oct. 19 to 1952 Jan. 17 (taken at an average interval of ∼ 110 days) shows nothing at this location (limiting mag ∼ 11.5 in 1930 to ∼ 13.5 in 1950).

R. M. Wagner, LBT Observatory; and J. P. Halpern and M. Jackson, Columbia University, communicate: “Visual examination with the Hiltner 2.4-m telescope (+ television acquisition system) at the MDM Observatory on Kitt Peak on Jan. 9.27 UT reveals the presence of a new or brightened stellar object (mag ∼ 10) near the position reported by Brown. When compared with the Digitized Sky Survey, the object appears virtually coincident with a USNO-A2.0 star (red mag 14.3) located at α = 7^{h}04^{m}04^{s}.846, δ = −3^{o}50′51′′.12 (equinox 2000.0), although the DSS image is a blend of at least two distinct stellar objects. A spectrum (range 393–856 nm, resolution 0.5 nm) reveals numerous absorption lines and four prominent emission lines exhibiting P-Cyg profiles. Three of these can be identified with Ba II (multiplet 2) at 585.5, 614.2, and 649.7 nm. In addition, strong Na I 589.7-nm and weak Hα emission (equivalent width 0.1 nm, FWHM 200 km/s) are present and exhibit P-Cyg profiles. Other features include many absorption lines of Fe II in the region 500–560 nm, O I 777.4-nm, and Ca II at 859.8 and 854.2 nm. However, the spectrum is not typical of Fe- or He/N-type classical novae early in their evolution.”

CCD broadband-red magnitudes with an RG-1 filter (cf. McCullough and Thakkar 1997, PASP 109, 1264) obtained by J. Bedient, Honolulu, HI: 2001 Dec. 22.284, 24.279, and 26.273 UT, [12; 2002 Jan. 1.257, 9.64; 2.254, 9.34; 3.251, 9.10; 4.249, 9.05; 5.246, 8.80; 8.238, 9.02; 9.235, 8.88.

2002 January 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green